# Random regular graph states are complex at almost any depth

### v1

@@ 1.2.3 Universality of regular graphs of intermediate degree
Let $G$ be a random $d$–regular graph on $n$ vertices, with $d=n^{c}$ where $10.5<c<$ 1. Then, with probability $1-o(1)$, it contains a square grid graph on $v$ vertices, for any $v=o(n^{k})$, with $d=\mathsf{min}\left\{\frac{1-c}{2},\frac{c}{3}\right\}$, as an induced subgraph.
This theorem is proven using the switching technique due to McKay and Wormald [52]. This is a standard technique in random regular graph theory, but is usually applied in situations where the subgraph to be found is of constant size (whereas for us it must grow reasonably fast with $n$). Because of this, and because these techniques are not widespread in the quantum computing literature, we give explicit switching calculations in Section 4.1.
A limitation of this result is the fact that $c>0.5$. This seems to be an unavoidable fact, as the expected number of grid graphs goes to zero whenever $c<0.5$. This threshold behavior is observed even for grid graphs of constant size [46]. We can get around this limitation by considering sparsified grid graphs, constructing by replacing every edge in an $L\times L$ grid graph by $L$ edges and $L-1$ vertices in a line. The resulting graphs are very line-like asymptotically yet are still universal (with a polynomial space overhead): we can recover the $L\times L$ grid graph as a vertex minor by measuring all the added vertices in the $Y$ basis. For these graphs we can prove the stronger statement:
@@
Let $G$ be a random $d$–regular graph on $n$ vertices, with $d=n^{c}$ where $10.5<c<$ 1. Then, with probability $1-o(1)$, it contains a sparsified square grid graph on $v$ vertices, for any $v=o(n^{k})$, with $d=\mathsf{min}\left\{\frac{1-c}{2},\frac{c}{3}\right\}$, as an induced subgraph.
Corollary 31 implies that random $n^{c}$-regular graphs of intermediate degree $c\in(0,1)$ are universal resources for measurement-based quantum computing when measured in an arbitrary product basis. Note, however, that finding the induced sparsified grid graph may be (and probably is) a computationally difficult problem. Therefore, the standard reduction from sampling to computing probabilities of post-selected BQP computations may not be possible in polynomial time. However, since identifying a grid graph can be done with access to an NP oracle, the reduction is possible in the polynomial hierarchy. Thus, while an efficient exact sampling algorithm in an arbitrary basis from a grid graph collapses the polynomial hierarchy to the third level [33], it still collapses it to the fourth level for graphs containing induced sparsified grid graphs.
@@ 1.2.4 Absence of a geometric entanglement barrier for uniformly random graph states
$E_{g}(\ket{\psi})=-\log\left(\max_{\alpha\in\mathrm{PROD}_{n}}|\innerproduct{\alpha}{\psi}|^{2}\right).$ (3)
where $\mathrm{PROD}_{n}$ is the set of product states, then any NP problem solved by an MBQC with |$\ket{\psi}$ as a resource state can be solved by a classical computer (with access to randomness) in time $O\big{(}\mathrm{poly}(n)2^{n-E_{g}(\psi)}\big{)}$. In light of the failure of the arguments above in the $d=\Theta(n)$ regime one can wonder whether this entanglement barrier shows up for random graph states (of degree $cn$ with $c\in(0,1)$). We give evidence that this is not the case by proving an upper bound on the geometric entanglement of uniformly random graph states.
@@
Choose a graph state |$\ket{G}$ uniformly at random. There exists constants $c,C$ such that
$\mathbb{P}\bigg{[}E_{g}(\ket{G})\geq n-cn^{1/4}/\log(n)\bigg{]}\leq C.$ (4)
@@
Choose a stabilizer state |$\ket{S}$ uniformly at random. There exists a constant $c$ such that
$\mathbb{P}\bigg{[}E_{g}(\ket{S})\leq n-c\sqrt{n}\log(n)\bigg{]}\leq O(2^{-\sqrt{n}}).$ (5)
This theorem is a straightforward application of the representation theory of the Clifford group. Because a constant fraction of stabilizer states are graph states (up to phases) this leads to a similar statement for graph states.
@@
Choose a graph state |$\ket{G}$ uniformly at random. There exists a constant $c$ such that
$\mathbb{P}\bigg{[}E_{g}(\ket{G})\leq n-c\sqrt{n}\log(n)\bigg{]}\leq O(2^{-\sqrt{n}}).$ (6)
@@ 1.3.3 High degree graph states
The main results on graph states of high degree are LABEL:thm:geom_ent_upper and LABEL:thm:geom_ent_lower_stab, providing lower and upper bounds on the geometric entanglement. LABEL:thm:geom_ent_lower_stab follows a relatively standard path, approximating the continuous optimization in the geometric entanglement by a discrete one through an epsilon net, followed by a union bound and a tail bound on the overlaps with fixed product states. There is some subtlety in that we require a rather small epsilon-net for the union bound to be non-trivial. To that end we extend a nice trick from random matrix theory to the geometric entanglement in Lemma 35, proving that we can obtain a multiplicative approximation to the maximum in the geometric entanglement using a relatively small epsilon net. The subsequent tail bound is then provided by the moments of random stabilizer states, developed in [27]. Some care must be taken to choose the right moment here, as the moments of random stabilizer states grow too fast for a straightforward exponential generating function approach to work [41].
@@ 1.4 Discussion and outlook
Our results are the first to explore the average-case complexity of random graph states with bounded degree. They give evidence towards the average-case complexity of simulating graph states of any degree $22<d\lesssim n/$ 2 using different types of results relating to the hardness and simulability of random graph states. In particular, our results interpolate between these extremely degree regimes. This interpolation runs via the degree of regular graphs and thus extends the results of [23] to the average case.
@@ 1.4.1 Constant degree graph states and random matching IQP circuits
Similar constructions of universal random circuits based on random matchings have previously been studied numerically and experimentally in [12]. Since commuting IQP circuits have often been precursors to results for random circuits (e.g. in terms of complexity [7, 6] and noisy simulation [8, 22, 2]), our result may thus also help in the rigorous study of constant-depth random circuits in arbitrary geometries. In particular, it is worth considering our results in relation to the results of [54], who give evidence that average-case hardness fails for constant-depth Haar-random circuits, as well as to the results of [12] who give evidence that the simulation complexity of random circuits on different connectivities remains bounded at very low depths. To see why our results are different, we observe that both of those works consider random circuits in which arbitrary single-qubit rotation gates can be applied throughout the circuit. It may indeed be that in this case, average-case hardness requires a strictly super-constant circuit depth. In fact, at least depth $\Omega(\log\log n)$ is required for anticoncentration in models that are invariant under Haar-random single qubit gates [65]. Our results circumvent this lower bound by restricting the local bases we measure in to the $X$-$Y$ plane. Hence, the fact that we only measure in the $X$-$Y$ plane is critical to the anticoncentration in constant depth we show. It remains an interesting question, however, how the simulation complexity of such random circuits depends on the circuit depth. As a first step towards this, it would be interesting to consider the random matching IQP circuits with measurements in a Haar-random single qubit basis. By the result of [65], anticoncentration will fail for constant depth, and thus this model would be helpful to understand the mechanism governing the arising lower bound.
@@ 1.4.2 Intermediate degree graph states and universality for measurement-based computation
[…]
@@ 1.4.3 High degree graph states
In the regime of high degree we also lack a full characterization of the complexity of graph states, in particular the question of classical simulability is not fully settled. While we have excluded the simulation algorithm given in [24], we can not exclude the existence of more tailor-made classical algorithms for MBQC with random linear-degree graph states. In Section 5.3 we described a simulator where all Pauli measurements are explicitly calculated and only non-Pauli measurement are simulated. We believe that even this more sophisticated algorithm fails, but can not prove it as of yet. We leave this as a conjecture. It would also be interesting to prove variants of our results for random $cn$-regular graph states with $20c<1/$ 2, which we believe can be done through the sandwich theorems from random graph theory [25].
@@ Acknowledgements
JH would like to thank Ingo Roth and Jop Briet for useful suggestions on the proof of LABEL:thm:geom_ent_upper, and Michael Walter and Marcel Hinsche for helpful discussions on the Clifford group. DH thanks Michael Gullans for many inspiring discussions on stat-mech mappings and IQP circuits. We also thank Daniel Grier, Jackson Morris, and Debbie Leung for discussions on vertex minors and simulation algorithms. We gratefully acknowledge the hospitality of the Simons Institute for the Theory of Computing in the spring of 2024, supported by DOE QSA grant #FP00010905, where part of this work was conducted. DH acknowledges funding from the US DoD through a QuICS Hartree Fellowship and from the Simons Institute for the Theory of Computing, supported by DOE QSA grant #FP00010905. JH acknowledges funding from the Dutch Research Council (NWO) through a Veni grant (grant No.VI.Veni.222.331) and the Quantum Software Consortium (NWO Zwaartekracht Grant No.024.003.037).
@@ 2.1 Graphs, graph states and stabilizer states
$m(G)=\text{max}\bigg{\{}\frac{|E|_{H^{\prime}}}{|V|_{H^{\prime}}},\leavevmode\nobreak\ H^{\prime}\subseteq G,\leavevmode\nobreak\ \leavevmode\nobreak\ |V|_{H^{\prime}}>0\bigg{\}},$ (7)
An $n$-qubit graph state |$\ket{G}$ is defined in terms of an $n$-vertex graph $G$ as
|$\ket{G}=\prod_{(i,j):U_{G}[i,j]=1}CZ_{i,j}\ket{+^{n}}=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}(-1)^{x^{T}U_{G}x}\ket{x},$ (8)
where $U_{G}$ is the upper triangle of the adjacency matrix $A_{G}$ of $G$ and the inner product is taken over $\mathbb{F}_{2}$. Graph states are a type of stabilizer state (which we will denote by |$\ket{S}$), which means they are the joint $+1$ eigenvectors of a set of $2^{n}$ mutually commuting Pauli matrices. There are $2^{n(n-1)/2}$ graph states and $2^{n}\prod_{i=1}^{n}(2^{i}+1)$ stabilizer states. A key property of stabilizer states is the following expectation value (taken uniformly over the stabilizer states), first derived in [27]:
$\mathbb{E}_{\ket{S}}\ket{S}\!\bra{S}^{\otimes t}=\frac{1}{2^{n}\prod_{i=0}^{t-2}(2^{n}+2^{i})}\sum_{T\in\Sigma_{t,t}}R(T),$ (9)
@@ 2.2 Random graph models
$G_{p}(d,n)$ for $dn$ even is the distribution over regular multigraphs of degree $d$ generated by choosing a uniformly random matching $M$ on the set $\{(i,j)\;\|\;i\in[n],j\in[d]\}$ and adding an edge between $i,\hat{i}$ whenever $(i,j),(\hat{i},\hat{j})$ is an edge in $M$.
@@ 2.3 Krawtchouck polynomials
Krawtchouk polynomials are a family of polynomials that prominently appear in classical error correction codes and boolean analysis. We will need them in the proof of Theorem 17. We will briefly recap their definition here as well as a number of upper bounds. The (binary) Krawtchouk polynomial of degree $i$ and size $N$ is defined as
@@
This is a rather straightforward upper bound but it will service almost all of our needs (it is also surprisingly close to being tight, see [45]). However to cover certain parameter regimes we will also need a more sophisticated bound from [15, Corollary 16] of the form

### v2

@@ 1.2.3 Universality of regular graphs of intermediate degree
Let $G$ be a random dd–regular graph on $n$ vertices, with $d=n^{c}$ where $0.5<c<1$. Then, with probability 1o(1)1-o(1), it contains a square grid graph on $v$ vertices, for any $v=o(n^{k})$, with $k=\mathsf{min}\left\{\frac{1-c}{2},\frac{c}{3}\right\}$, as an induced subgraph.
This theorem is proven using the switching technique due to McKay and Wormald [71]. This is a standard technique in random regular graph theory, but is usually applied in situations where the subgraph to be found is of constant size (whereas for us it must grow reasonably fast with $n$). Because of this, and because these techniques are not widespread in the quantum computing literature, we give explicit switching calculations in Section˜4.1.
A limitation of this result is the fact that $c>0.5$. This seems to be an unavoidable fact, as the expected number of grid graphs goes to zero whenever $c<0.5$. This threshold behavior is observed even for grid graphs of constant size [63]. We can get around this limitation by considering sparsified grid graphs, constructing by replacing every edge in an $L\times L$ grid graph by $L$ edges and L1L-1 vertices in a line. The resulting graphs are very line-like asymptotically yet are still universal (with a polynomial space overhead): we can recover the $L\times L$ grid graph as a vertex minor by measuring all the added vertices in the $Y$ basis. For these graphs we can prove the stronger statement:
@@
Let $G$ be a random dd–regular graph on $n$ vertices, with $d=n^{c}$ where $0.5<c<1$. Then, with probability 1o(1)1-o(1), it contains a sparsified square grid graph on $v$ vertices, for any $v=o(n^{k})$, with $k=\mathsf{min}\left\{\frac{1-c}{2},\frac{c}{3}\right\}$, as an induced subgraph.
Corollary˜31 implies that random $n^{c}$-regular graphs of intermediate degree $c\in(0,1)$ are universal resources for measurement-based quantum computing when measured in an arbitrary product basis. Note, however, that finding the induced sparsified grid graph may be (and probably is) a computationally difficult problem. Therefore, the standard reduction from sampling to computing probabilities of post-selected BQP computations may not be possible in polynomial time. However, since identifying a grid graph can be done with access to an NP oracle, the reduction is possible in the polynomial hierarchy. Thus, while an efficient exact sampling algorithm in an arbitrary basis from a grid graph collapses the polynomial hierarchy to the third level [48], it still collapses it to the fourth level for graphs containing induced sparsified grid graphs.
@@ 1.2.4 Absence of a geometric entanglement barrier for uniformly random graph states
[…]
@@
[…]
@@
[…]
This theorem is a straightforward application of the representation theory of the Clifford group. Because a constant fraction of stabilizer states are graph states (up to linear phases, which leave the geometric entanglement unchanged) this leads to a similar statement for graph states.
@@
[…]
@@ 1.3.3 High degree graph states
The main results on graph states of high degree are LABEL:thm:geom_ent_upper and LABEL:thm:geom_ent_lower_stab, providing lower and upper bounds on the geometric entanglement. LABEL:thm:geom_ent_lower_stab follows a relatively standard path, approximating the continuous optimization in the geometric entanglement by a discrete one through an epsilon net, followed by a union bound and a tail bound on the overlaps with fixed product states. There is some subtlety in that we require a rather small epsilon-net for the union bound to be non-trivial. To that end we extend a nice trick from random matrix theory to the geometric entanglement in Lemma˜35, proving that we can obtain a multiplicative approximation to the maximum in the geometric entanglement using a relatively small epsilon net. The subsequent tail bound is then provided by the moments of random stabilizer states, developed in [41]. Some care must be taken to choose the right moment here, as the moments of random stabilizer states grow too fast for a straightforward exponential generating function approach to work [56].
@@
1.4 Context and prior work
Our work builds on first steps made in Ref. [37] that classify the simulation complexity and entanglement properties in the worst case over the choice of $d$-regular graphs. There, it was shown that as $d$ is increased, the simulation complexity and entanglement properties undergo two phase transitions, providing a tight link between complexity and entanglement: For $d\leq 2$ and $d\geq n-3$ the entanglement is low and simulating single-qubit measurements in any basis is classically easy for all $d$-regular graph, while for any other value of $d$ there is a $d$-regular graph for which simulations are classically intractable and the multipartite entanglement is high. But are these hard instances isolated in their respective regularity class or are most instances of a class hard? In many cases, average-case complexity significantly differs from worst-case complexity, most famously so for $\mathsf{NP}$-problems, where often most instances are in fact efficiently solvable.
Thus, our results on average-case hardness of random $d$-regular graph states significantly strengthen the connection between the connectivity of a graph and the entanglement properties of the corresponding state and provide complexity-theoretic evidence that classical intractability is a generic feature of multipartite entangled states. Generally speaking, though, there are only few techniques to address average-case complexity such as random self-reducibility—these techniques primarily involve reducing the average-case problem to proving average-case hardness of computing the permanent of a matrix, for e.g. see [88, 1, 11]—which work for Haar random gate sets since they are continuously parameterized but are not applicable to the discrete randomness of random graphs. This is why the average-case complexity of random graph states is a qualitatively different question from both their worst-case complexity, and average-case complexity of continuous ensembles.
From a different perspective, our work can be viewed as a first attempt to answer the question whether random regular graph states are resources for measurement-based computing (MBQC) (and in the regime of $d=\Theta(n^{c})$ we answer in the affirmative). Assessing which graphs form resource states for MBQC has a rich history [17, 76, 47, 92, 5] and has motivated the idea of computational phases of matter [19, 18, 65, 20, 34, 32, 83]. This question has also arisen when studying the impact of noise on computational power. Concretely, Browne et al. [8] studied the impact of erasure noise on MBQC on grid graphs in terms of percolation phenomena. Here, at each lattice site, there is a finite probability that the qubit on that lattice site is erased, resulting in a random cluster state as originally introduced by Briegel and Raussendorf [17]. Below a certain value of that probability—known as the percolation threshold—noisy grid graphs can then be classically simulated, whereas above that probability, noisy grid graphs remain universal resource states. Equivalently, we can view this setting as the random graph state ensemble given by random subgraphs of the grid graph [17], which is a rather different ensemble than the ones studied in this paper.
@@ 1.5 Discussion and outlook
Our results are the first to explore the average-case complexity of random graph states with bounded degree. They give evidence towards the average-case complexity of simulating graph states of any degree $2<d\lesssim n/2$ using different types of results relating to the hardness and simulability of random graph states. In particular, our results interpolate between these extremely degree regimes. This interpolation runs via the degree of regular graphs and thus extends the results of [37] to the average case.
@@ 1.5.1 Constant degree graph states and random matching IQP circuits
Similar constructions of universal random circuits based on regular graphs with perfect matchings have previously been studied numerically and experimentally in [24]. Since commuting IQP circuits have often been precursors to results for random circuits (e.g. in terms of complexity [14, 13] and noisy simulation [15, 36, 4]), our result may thus also help in the rigorous study of constant-depth random circuits in arbitrary geometries. In particular, it is worth considering our results in relation to the results of [73], who give evidence that average-case hardness fails for constant-depth Haar-random circuits, as well as to the results of [24] who give evidence that the simulation complexity of random circuits on different connectivities remains bounded at very low depths. To see why our results are different, we observe that both of those works consider random circuits in which arbitrary single-qubit rotation gates can be applied throughout the circuit. It may indeed be that in this case, average-case hardness requires a strictly super-constant circuit depth. In fact, at least depth $\Omega(\log\log n)$ is required for anticoncentration in models that are invariant under Haar-random single qubit gates [84]. Our results circumvent this lower bound by restricting the local bases we measure in to the $X$-$Y$ plane. Hence, the fact that we only measure in the $X$-$Y$ plane is critical to the anticoncentration in constant depth we show. It remains an interesting question, however, how the simulation complexity of such random circuits depends on the circuit depth. As a first step towards this, it would be interesting to consider the random matching IQP circuits with measurements in a Haar-random single qubit basis. By the result of [84], anticoncentration will fail for constant depth, and thus this model would be helpful to understand the mechanism governing the arising lower bound.
@@ 1.5.2 Intermediate degree graph states and universality for measurement-based computation
Our results focus on the question of whether random graph states can be classically simulated, which is a priori a question which is distinct from the question whether random graph states are universal resources for measurement-based quantum computation. While some of our results (in the intermediate regime) explicitly make use of universal resources, it is interesting to ask in what sense or to what extent random graph states in different regimes can function as universal resource states. To answer this question, it must first be made clear that the notion of a “universal resource” can have very different meanings. A natural notion of universality is one in which we just ask that for any number of qubits $n$, the family of states contains a state on $m\geq n$ qubits such that using measurements and feed-forward an arbitrary unitary can be implemented on $n$ of the $m$ qubits up to some precision threshold [29]. Note that this notion is not concerned with efficiency, and hence the graphs containing a 2D grid graph in the intermediate regularity regime are certainly examples of universal resources in this sense. The universal resource is even efficient in the sense that $m={\mathsf{poly}}(n)$. However, note that these graphs cannot be efficiently found in general, and finding them may require the solution of an NP problem (showing this is an interesting open question). In order to computationally exploit a universal resource, it is a prerequisite that a sufficiently large 2D subgraph can also be efficienlty found, since the only known efficient constructions make use of those graphs. This raises the question whether in the intermediate degree regime certain random ensembles contain grid graphs as subgraphs that can also be efficiently found.
@@ 1.5.3 High degree graph states
In the regime of high degree we also lack a full characterization of the complexity of graph states, in particular the question of classical simulability is not fully settled. While we have excluded the simulation algorithm given in [38], we can not exclude the existence of more tailor-made classical algorithms for MBQC with random linear-degree graph states. In Section˜5.3 we described a simulator where all Pauli measurements are explicitly calculated and only non-Pauli measurement are simulated. We believe that even this more sophisticated algorithm fails, but can not prove it as of yet. We leave this as a conjecture. It would also be interesting to prove variants of our results for random $cn$-regular graph states with $0c<1/2$, which we believe can be done through the sandwich theorems from random graph theory [39].
1.5.4 Towards analyzing noisy, architecture-constrained graph states
Aside from the different regimes of the regularity parameter just discussed, an interesting open question raised by our work is the complexity of graph states that can be naturally realized in different quantum computing architectures. The arbitrary connectivity required to generate the random regular graphs may be difficult to realize in practice. For instance, while in principle reconfigurable atom or ion arrays [9, 24] arbitrary connectivity can be realized, there may be more natural random ensembles of graph states, for instance, subgraphs of a high-dimensional hypercube [51]. In other architectures, such as superconducting qubits, even more constrained low-dimensional lattice geometries are imposed [3].
A further interesting question in the context of more realistic circuit ensembles is to what extent noise affects the output distribution. In the literature on sampling from the output distributions of random quantum circuits, a prominent quantity is the cross-entropy benchmark (XEB) score, which generalizes anticoncentration to noisy circuits and serves as a measure of quality of the sampled distribution [13, 3, 59]. An important result in this context has been that for local noise rates on the order of $\lesssim 1/n$ and where the ideal output distribution anticoncentrates, the XEB can be used as a proxy of the many-body fidelity of the pre-measurement state [22, 70, 93], and moreover, that this is the regime in which there are no exploits that classical algorithms might use to achieve a high XEB score significantly more efficiently than brute-force simulations. In contrast, in the regime where the noise rate is larger than $1/n$, such exploits exist [40]. In fact, as the depth is increased Similar statements can also be made for IQP circuits [51, 9], and it is an interesting open question to analyze the XEB for noisy random graph state ensembles.
@@ Acknowledgements
JH would like to thank Ingo Roth and Jop Briet for useful suggestions on the proof of LABEL:thm:geom_ent_upper, and Michael Walter and Marcel Hinsche for helpful discussions on the Clifford group. DH thanks Michael Gullans for many inspiring discussions on stat-mech mappings and IQP circuits. We also thank Daniel Grier, Jackson Morris, and Debbie Leung for discussions on vertex minors and simulation algorithms. We gratefully acknowledge the hospitality of the Simons Institute for the Theory of Computing in the spring of 2024, supported by DOE QSA grant #FP00010905, where part of this work was conducted. DH acknowledges funding from the Simons Institute for the Theory of Computing, supported by DOE QSA. JH acknowledges funding from the Dutch Research Council (NWO) through a Veni grant (grant No.VI.Veni.222.331) and the Quantum Software Consortium (NWO Zwaartekracht Grant No.024.003.037).
@@ 2.1 Graphs, graph states and stabilizer states
We begin by reviewing some standard graph theoretic notions. A graph $G$ is a set of vertices $V$ (usually the set $[n]$, and a set of vertices $E\subset V\times V$ connecting them). We denote by $\overline{G}$ the complement graph, which has the same vertex set, and the complementary set $\overline{E}\in V\times V$ of edges. We will occasionally be somewhat sloppy in notation and write $e\in G$ ($e\not\in G$)to indicate that $e$ is (not) an edge in $G$ (and thus $e\in E$ ($e\not\in E$)). Similarly we will sometimes write subsets of edges as $S\subseteq G$. A subgraph $H$ of $G$ is a obtained from $G$ by considering a subset of the edges $E^{\prime}\subseteq E$ and an induced subgraph $H$ is obtained by considering a subset of the vertices $V^{\prime}\subseteq V$, which has edges $E^{\prime}=E\cup V^{\prime}\times V^{\prime}$.
@@
$m(G)=\text{max}\bigg\{\frac{|E|_{H^{\prime}}}{|V|_{H^{\prime}}},\penalty 10000\ H^{\prime}\subseteq G,\penalty 10000\ \penalty 10000\ |V|_{H^{\prime}}>0\bigg\},$ (7)
[…]
@@ 2.2 Random graph models
[…]
@@ 2.3 Krawtchouck polynomials
Krawtchouk polynomials are a family of polynomials that prominently appear in classical error correction codes and boolean analysis. We will need them in the proof of Theorem˜17. We will briefly recap their definition here as well as a number of upper bounds. The (binary) Krawtchouk polynomial of degree $i$ and size $N$ is defined as
@@
This is a rather straightforward upper bound but it will service almost all of our needs (it is also surprisingly close to being tight, see [62]). However to cover certain parameter regimes we will also need a more sophisticated bound from [27, Corollary 16] of the form